1.     Suppose a consumer has income of 1000, and buys pizza and coke. The price of pizza is 100 and price of coke is 50. 

(i)    Draw a budget line that represents the set of bundles this individual can afford if they use all their income. (Put movies on the X axis). Label the places where the budget line intercepts each axis and the slope of the line.

(ii)  Suppose that in consumer equilibrium this individual consumes 7 pizza and 6 coke. What rules must hold in consumer equilibrium? Label this bundle in your drawing (call it point “A”). 

(iii)Draw an indifference curve associated with this bundle, and explain how its slope at the equilibrium point relates to the slope of the budget line.


1
Expert's answer
2021-05-12T07:26:26-0400

(i) A budget line in this case is a downward-sloping line which intercepts Y (coke) axis at the point (0; 20) and X (pizza) axis at the point (10; 0). The slope of the line is "k = (20 - 0)\/(0 - 10) = -2."

(ii) 7 pizza and 6 coke is on the budget line, because 100*7 + 50*6 = 1000.

(iii) An indifference curve associated with this bundle will pass through the point A, and will have almost the same slope as the budget line.